### v2

# Spatial super-resolution in nanosensing with blinking emitters

Abstract
We propose a method of spatial resolution enhancement in metrology (thermometry, magnetometry, pH estimation, and similar methods) with blinking fluorescent nanosensors by combining sensing with super-resolution optical fluctuation imaging (SOFI). To demonstrate efficiency of this approach, both numerical simulations and a model experiment with laser diodes, which model fluctuating nanoemitters, and intentional blurring of the image are performed. The 2nd, 3rd, and 4th order cumulant images provide improvement of the contrast and enable successful reconstruction of smaller features of the modeled temperature (or any other physical parameter) distribution relatively to the intensity-based approach. We believe that blinking fluorescent sensing agents being complemented with the developed image analysis technique could be utilized routinely in the life science sector for recognizing the local changes in the spectral response of blinking fluorophores, e.g. delivered targetly to the wanted cell or even organelle. It is extremely useful for the local measurements of living cells’ physical parameters changes due to applying any external “forces”, including disease effect, aging, healing, or response to the treatment.
nanosensors, quantum dots, fluorescence, super-resolution optical fluctuation imaging, spatial resolution, sub-Rayleigh imaging methods
I Introduction
Fluorescent nanosensors represent a powerful tool for measurement of temperature, magnetic field, pH, and other parameters with high accuracy at small scales [1, 2, 3, 4, 5, 6]. In particular, the dependence of their fluorescent signal on the mentioned parameters can be utilized for investigation of processes in living cells [2, 7, 8]. To ensure high spatial resolution, one needs to localize the nanoemitters precisely and to distinguish signals from closely located nanosensors. This requirement can be fulfilled, for example, by placement of fluorescent nanodiomonds or quantum dots at a manipulator needle [9, 10, 11, 12, 13]. This approach requires precise mechanical control and does not allow to analyze environment parameters at several points of the sample simultaneously. For a large number of nanosensors, simultaneously attached to the sample, one can apply deterministic super-resolution approaches based on selective (localized) excitation of emitters (stimulated emission depletion (STED), ground state depletion (GSD), reversible saturable optical fluorescence transitions between spin levels (spin-RESOLFT), structured illumination imaging) [14, 15, 16, 17, 18], as well as stochastic methods (stochastic optical reconstruction microscopy (STORM), photoactivated localization microscopy (PALM), points accumulation for imaging in nanoscale topography (PAINT)) [19, 20, 21, 22], ensuring independent localization of single emitters due to their blinking. Those approaches are also quite sophisticated and require the nanosensors to remain in the inactive state most of the time, thus increasing needed exposition time and potential photo-toxicity.
Alternatively to the super-resolution approaches, based on scanning or selective excitation (activation) of emitters, wide-field microscopy methods exploiting specific temporal statistics of fluorescent objects, i.e. stochastic blinking (in super-resolution optical fluctuation imaging — SOFI) [23, 24, 25] or photon antibunching [26, 27, 28] have been successfully implemented. The idea is based on postprocessing of time-dependent fluorescent signal: its efficient exponentiation by considering correlations instead of raw intensity, elimination of cross-terms from different emitters, and subsequent enhancement of spatial resolution. Analysis of the fluorescence blinking does not prevent usage of the “degrees of freedom” suitable for the environment sensing, such as the dependence of the spectrum or fluorescence lifetime on the temperature, acidity, pressure, etc. In the current contribution we suggest to combine sensing capabilities of nanoparticles with SOFI-based super-resolution for more accurate localization of the signal source. Is is important to note that the information, used by SOFI to achieve sub-Rayleigh spatial resolution, is already present in the collected fluorescent signal and does not introduce any additional complications to the measurement.
Combining SOFI with sensing is not as trivial as it may seem at the first glance. First, we focus on the case when the parameter of interest may vary at the scales smaller that the classical resolution limit, determined by the point spread function (PSF) of the used microscope. Therefore, the assumptions of slowly varying parameters, used for example in balanced SOFI (bSOFI) [29], are not generally valid here. Second, SOFI imposes less restrictions on separate frames of the acquired timeseries in comparison with single-molecule localization techniques [14, 15, 16, 17]: it allows smaller numbers of detected photons per frame and higher probability of closely positioned emitters being active during the same frame. Thus, one cannot reliably perform reconstruction of the sensed environmental parameter for a single-frame dataset or for a separate localized emitter as it is done in spectrally resolved STORM, PAINT, and PALM techniques [30, 31, 32, 33, 34]. Third, SOFI requires analysis of the whole acquired timeseries, relies on specific temporal fluctuations of the signal, and yields nonlinear dependence of the response on the emitters’ brightness [29, 35, 36]. Therefore, conversion of each separate frame into the investigated parameter map before processing of the entire timeseries would alter the fluctuations and preclude further application of SOFI for spatial resolution enhancement. On the other hand, application of SOFI to multichannel frame data (for example, spectra) would tremendously enhance the shot noise, while nonlinearity of the response would require additional treatment.
To resolve the outlined issues and combine SOFI with sensing, we propose to “compress” each frame to a two-channel image, apply SOFI to the timeseries corresponding to each of the two channels separately, and infer the sensed parameter map from the obtained cumulant images for those channels. Linearity of the compression ensures preservation of the temporal fluctuations of the signal and applicability of SOFI to spatial resolution enhancement. The idea of using two channels for single-parameter inference is closely related to ratiometric imaging [37, 38], where two spectral channels can be used for self-referencing [39] to reduce the adverse effect of absolute intensity fluctuations [37, 39, 40]. Nonlinearity of the resulting response of SOFI [29, 35, 36] can be compensated analytically by inversion of the nonlinear relation between the sensed parameter of interest and cumulant values.
The idea of combined usage of blinking statistics of fluorophores for super-resolution and certain other degree of freedom for getting additional information is somewhat close to so called “color SOFI” [41, 42, 43, 44, 45] — super-resolution technique, which applies SOFI to several (typically two) spectral channels. However, the tasks, solved by those two approaches, are different. Color SOFI deals with discrimination between several (two or three) discrete types of fluorescent emitters, serving as biological labels. In contrast, the method proposed in this paper, deals with continuous-variable sensing: the emitters are of the same type, but their fluorescent signal differs due to variations of local environment.
To demonstrate principles of the proposed approach and show its feasibility, we first applied in to multiple synthetic (simulated) datasets and quantified the performance and resolution enhancement. Information from both auto- and cross-cumulants [24] with zero time lag was combined to reduce the influence of the shot noise and improve the accuracy of sensing. The mean square error of the sensed parameter estimation scaled as $1:0.28:0.0075$ for the 1st, 2nd, and 4th order cumulants for the specific considered model. To get closer to real-world testing of the approach in non-ideal conditions (without exact prior knowledge of PSF and with measurement imperfections), we also performed a model experiment mimicking blinking fluorescent nanoemitters. We used classical photodiods with the controlled signal imitating random switching between “bright” and “dark” states. Similarly to model (proof-of-principle) super-resolution experiments in quantum imaging [46, 47, 48, 49, 50], the image was intentionally blurred (in our case, by inaccurate focusing on the detection plane) so that the spots produced by neighbor “emitters” were strongly overlapping. Model “temperature” of the “emitters” was encoded in the “fluorescence lifetime” — the delay between the “excitation pulse” and the emission response. The subsequent application of the proposed SOFI-based sensing approach allowed to enhance spatial resolution approximately twice even for the simplest version of SOFI with zero time lag autocumulants.
The paper is organized as follows. In Section II, we discuss basic ideas of the approach and list the procedures constituting the proposed technique of spatially super-resolving sensing. Section III is devoted to application of the approach to synthetic datasets and quantification of the resolution (accuracy of parameter estimation) improvement. Section IV describes the model experiment, data acquisition, and reconstruction results for the model “temperature” distribution. Conclusions on the results and possible directions of further research are presented in Section IV.5.
II Basic ideas
II.1 Super-resolution optical fluctuations imaging
The commonly used model of wide-field fluorescence imaging includes a sample composed of $N$ incoherent florescent emitters, located at positions $\textbf{r}_{k}$ [23]. Fluctuations-based approaches assume that the emitters are independently blinking, stochastically switching between “bright” and “dark” states. The acquisition time is split into a finite number of frames. The fluorescence signal in the frame $t$ ($t=1$, …, $N_{\text{f}}$, where $N_{\text{f}}$ is the number of frames) can be described as [23]
$F(\mathbf{r},t)=\sum_{k}U(\mathbf{r}-\mathbf{r}_{k})\varepsilon_{k}s_{k}(t),$ (1)
where $U(\Delta\mathbf{r})$ is the point-spread function (PSF) of the optical system, $\varepsilon_{k}$ is the constant molecular brightness of the $k$-th emitter, and $s_{k}(t)$ is its time-dependent fluctuation ($s_{k}(t)=1$ if the $k$-th emitters remains “bright” during the $t$-th frame; $s_{k}(t)=0$ for the “dark” state).
The idea of resolution enhancement in the wide-field microscopy (including quantum imaging, SOFI, antibunching imaging) is to construct signal with power-law dependence on $U(\Delta\mathbf{r})$ and to take into account that, for a localized PSF, $U^{n}$ is approximately $\sqrt{n}$ times narrower than $U$ itself. For a single emitter ($N=1$), just raising the signal $F(\mathbf{r},t)$ to certain power $n>1$ would be sufficient for achieving the effect. For $N>1$ emitters, exponentiation of the signal leads to additional cross-terms, hindering resolution of the emitters. For example, for $n=2$ the squared signal equals
$F^{2}(\mathbf{r},t)=\sum_{k}U^{2}(\mathbf{r}-\mathbf{r}_{k})\varepsilon_{k}^{2}s_{k}^{2}(t)\\ +\sum_{k,k^{\prime}:k^{\prime}\neq k}U(\mathbf{r}-\mathbf{r}_{k})U(\mathbf{r}-\mathbf{r}_{k^{\prime}})\varepsilon_{k}\varepsilon_{k^{\prime}}s_{k}(t)s_{k^{\prime}}(t).$ (2)
To ensure resolution enhancement, one needs to construct a polynomial function of the fluorescence signal with excluded cross-terms [23]. For example, variance of the signal considered as a time series of frames both is quadratic relatively to PSF and has zero cross-terms for independent blinking of the emitters:
$\operatorname{Var}_{t}[F(\mathbf{r},t)]=\sum_{k}U^{2}(\mathbf{r}-\mathbf{r}_{k})\varepsilon_{k}^{2}\operatorname{Var}_{t}[s_{k}(t)].$ (3)
The approach can be extended to higher-order polynomials by using cumulants instead of variance [23]:
$C_{t}^{(n)}[F(\mathbf{r},t)]=\sum_{k}U^{n}(\mathbf{r}-\mathbf{r}_{k})\varepsilon_{k}^{n}C_{t}^{(n)}[s_{k}(t)].$ (4)
The expression stems from linearity of cumulants relatively to a sum of independent random variables: $C^{(n)}[X_{1}+X_{2}]=C^{(n)}[X_{1}]+C^{(n)}[X_{2}]$ for independent $X_{1}$ and $X_{2}$.
Several lower-order cumulants for a random variable (process) $f(t)$ are defined as
$C_{t}^{(1)}[f(t)]=\operatorname{E}_{t}[f(t)],$ (5)
$C_{t}^{(2)}[f(t)]=\operatorname{E}_{t}\bigl{[}\bigl{(}f(t)-\operatorname{E}_{t}[f(t)]\bigr{)}^{2}\bigr{]},$ (6)
$C_{t}^{(3)}[f(t)]=\operatorname{E}_{t}\bigl{[}\bigl{(}f(t)-\operatorname{E}_{t}[f(t)]\bigr{)}^{3}\bigr{]},$ (7)
$C_{t}^{(4)}[f(t)]=\operatorname{E}_{t}\bigl{[}\bigl{(}f(t)-\operatorname{E}_{t}[f(t)]\bigr{)}^{4}\bigr{]}-3\bigl{(}C_{t}^{(2)}[f(t)]\bigr{)}^{2},$ (8)
and correspond to variance, asymmetry, and excess kurtosis respectively. Here, $\operatorname{E}_{t}[f(t)]$ denotes time averaging of the random values $f(t)$ (defined at each frame $t$) over an infinite number of frames $N_{\text{f}}$. The random process of blinking is assumed to be stationary (with negligible effect of bleaching) so that ergodicity can be applied to the definition of cumulants. For finite acquisition time, expectation values are estimated as finite-sample averages:
$\operatorname{E}_{t}[f(t)]\approx\frac{1}{N_{\text{f}}}\sum_{t=1}^{N_{\text{f}}}f(t).$ (9)
Further, one can define also cross-cumulants and introduce non-zero time lags [24]:
$C_{t}^{(n)}[F(\mathbf{r}^{(1)},t+\tau_{1}),\ldots,F(\mathbf{r}^{(n)},t+\tau_{n})]\\ \propto\sum_{k}U^{n}(\bar{\mathbf{r}}-\mathbf{r}_{k})\varepsilon_{k}^{n}C_{t}^{(n)}[s_{k}(t+\tau_{1}),\ldots,s_{k}(t+\tau_{n})],$ (10)
where $\bar{\mathbf{r}}=\sum_{i=1}^{n}\mathbf{r}^{(i)}/n$.
To increase spatial resolution, SOFI deals with stochastic fluctuations of the total fluorescence signal, integrated over each frame duration and certain spectral range, rather than on spectral or temporal intensity distribution inside a frame. The technique itself does not impose any specific constraints on time or spectral windows, used during signal acquisition within frames. That property is crucial for the proposed sensing approach and was efficiently used in the color SOFI [41, 44]. It is worth noting, that there also exists an alternative approach of examining the emitters’ surrounding by analysis of the parameters of fluctuation already used for SOFI (for example, on-time ratio or blinking frequency) [29, 51, 52] rather than additional “degrees of freedom”.
II.2 Sensing with fluorescent nanoparticles
Sensing capabilities of fluorescent nanoparticles can be provided by dependence of their fluorescence lifetime, spectrum, or total intensity on the local parameters of the environment: temperature, pH, pressure, etc. [1, 2, 53, 54, 55, 56, 57, 58, 59, 60, 61] (Fig. 1). For example, carbon nanodots possess spectral dependence of fluorescence on pH and temperature [60] and can exhibit blinking required for SOfI [45]. Surface-induced charge-state switching of color centers in nanodiamonds [62, 63, 64, 65, 66, 67] enables combination of their vast sensing capabilities [1, 2, 53, 54, 55, 56, 57] with SOFI.
In other words, variation of the parameter of interest can change the total intensity of the fluorescent signal and (or) its distribution over certain parameter (variable) $x$ (wavelength, emission delay, etc.) within the frame. Variations of molecular surrounding [37, 39, 40] and blinking of the emitters (required for SOFI applicability) [29] introduce significant noise to the total (absolute) intensity of fluorescence. For that reason, further we focus on the latter case when the sensed environmental parameter influences the distribution of the fluorescent signal. Let $f(\mathbf{r},t,x)$ be the density of the signal over the variable $x$, connected with the total fluorescent signal $F(\mathbf{r},t)$ for the frame $t$ by the relation
[…]
For spectral response of the fluorophores to local changes of environment, $x$ may correspond to the emitted radiation wavelength and $f(\mathbf{r},t,x)$ be the spectral density of the signal. For sensing, based on fluorescence lifetime like in the case of SiV centers in diamond needles [58], $x$ is the delay between the excitation pulse and the detection of photons.
Figure 1: Sensing an environment parameter $\theta$ by its influence on fluorescence signal. The value of the parameter can be inferred from changes of the spectrum (a), fluorescence lifetime (b), or total intensity (c). The latter case is likely to yield worse sensing accuracy because of blinking-induced noise of absolute intensity [29]. The full spectral profile (a) can be compressed into two channels: with the optimal weights described in Appendix A (d) and with practical spectral windows having Gaussian profiles of sensitivity $Q_{1}(x)$ and $Q_{2}(x)$ (e). The channel $Y_{0}$ is defined as $Y_{1}+Y_{2}$ and corresponds to the weight $Q_{0}(x)=Q_{1}(x)+Q_{2}(x)$. The full temporal profile (b) can be compressed into two channels by using a simple threshold filter (f).
Eq. (1) can be rewritten in the following form for the density of the signal:
$f(\mathbf{r},t,x)=\sum_{k}U(\mathbf{r}-\mathbf{r}_{k})\varepsilon_{k}(x,\theta_{k})s_{k}(t),$ (12)
where the dependence on the $k$-th emitter intensity $\varepsilon_{k}(x,\theta)$ on the variable $x$ (wavelength or delay of the fluorescence signal) is affected by the local value of the environment parameter $\theta=\theta_{k}$ in the vicinity of the $k$-th emitter. To infer the local values $\theta(\mathbf{r})$ of the parameter in question, one needs to specify certain model for the dependence of $\varepsilon_{k}(x,\theta)$ on $\theta$. Typically, the emitters are assumed to be described by the same model $\varepsilon_{k}(x,\theta_{k})=\varepsilon_{k}\varepsilon(x,\theta_{k})$ up to a constant emitter-dependent brightness factor $\varepsilon_{k}$.
When trying to combine SOFI with sensing, it is tempting to apply cumulants construction to each value of $x$ (spectral or temporal channel) independently. However, as we discuss in the next section, splitting the fluorescent signal into a large number of channels leads to low photon numbers, significant shot noise, and inefficiency of high-order cumulants usage in SOFI. On the other hand, reconstruction of the whole distribution $\varepsilon(x,\theta(\mathbf{r}))$ is not needed to infer the spatial map of a single parameter $\theta(\mathbf{r})$. As we show here, the whole intensity profile can be “compressed” to just two scalar values with their ratio being sufficient for estimation of the parameter $\theta$. Accumulation of the intensity into two such channels effectively reduces the shot noise and restores applicability of SOFI for spatial super-resolution.
For now, let us consider sensing by a single probe nanoparticle localized at the position $\mathbf{r}_{0}$ and subjected to the environment with the parameter value $\theta$:
$f(\mathbf{r},t,x)=U(\mathbf{r}-\mathbf{r}_{0})\varepsilon_{0}(x,\theta)s_{0}(t).$ (13)
To eliminate additional inaccuracies, caused by blinking of the emitter, finite data acquisition time, or unknown brightness $\varepsilon_{0}$, one can work with normalized signal by dividing its density over the total value:
[…]
The idea of analyzing ratios instead of absolute values of the signal for reduction of fluctuations effects is widely used in ratiometric imaging [37, 38]. One can notice that the normalized signal depends on the normalized density of the emitter’s signal and is determined by the parameter of interest $\theta$, but does not include irrelevant geometrical information or the emitter brightness.
Instead of considering the whole normalized profile $\tilde{f}(\mathbf{r}_{0},t,x)$ for a specific frame $t$ and the emitter position $\mathbf{r}_{0}$, one can construct a quantity
[…]
for certain weight functions $Q_{0}(x)$ and $Q_{1}(x)$ and consider it as a “compressed” version of the measured signal (Fig. 1(d-f)). As shown in Appendix A, there exists the (locally) optimal choice of the weights (Fig. 1(d)), for which the quantity $Z$ contains exactly the same amount of information about the sensed parameter $\theta$ as the whole intensity profile $\tilde{f}(\mathbf{r}_{0},t,x)$ (i.e., represents sufficient statistics).
From practical point of view, Eq. (15) defines two detection channels $Y_{0}$ and $Y_{1}$ and proposes to use their ratio $Z$ for inference of $\theta$. In a real experiment, implementation of the optimal detection channels (with the optimal weight functions $Q_{0}(x)$ and $Q_{1}(x)$) can be complicated or impossible. Alternatively, one can just use the signals from two practically available channels $Y_{1}$ and $Y_{2}$ with their weight functions $Q_{1}(x)$ and $Q_{2}(x)$ (for example, corresponding to the signal accumulated over two spectral or temporal windows — Fig. 1(e,f)), define $Y_{0}=Y_{1}+Y_{2}$ (and $Q_{0}(x)=Q_{1}(x)+Q_{2}(x)$), and construct $Z$ according to Eq. (15). Strictly speaking, such measurement-induced choice of the weights is, generally, non-optimal and may lead to loss of information relatively to the full profile analysis.
Substituting Eq. (14) into Eq. (15), one can find the relation between the parameter $\theta$ and the compressed signal $Z$:
[…]
If the a priori constraints $\theta_{\text{min}}\leq\theta\leq\theta_{\text{max}}$ on the sensed $\theta$ ensure monotonicity of the dependence $\zeta(\theta)$, the function $\zeta(\theta)$ can be inverted and the estimator (map of the signal to the parameter of interest) can be constructed as
$\bar{\theta}=\zeta^{-1}(Z).$ (17)
For a particular case of linear dependence of the emitter’s signal density on the parameter of environment in the region of interest $[\theta_{\text{min}},\theta_{\text{max}}]$
$\varepsilon(x,\theta)=S_{0}(x)+S_{1}(x)\theta,$ (18)
the estimator can be constructed analytically:
$Z=\frac{A_{10}+A_{11}\theta}{A_{00}+A_{01}\theta}\quad\Rightarrow\quad\bar{\theta}=\frac{A_{00}Z-A_{10}}{A_{11}-A_{01}Z},$ (19)
where
[…]
It is worth mentioning that for the optimal weights $Q_{0}(x)$ and $Q_{1}(x)$, defined in Appendix A, we have $A_{ij}=\delta_{ij}$ and $\bar{\theta}=Z$.
II.3 Combining sensing and SOFI
The proposed idea of combining spatial resolution with sensing consists in applying SOFI separately to several channels relatively to the variable $x$, which defines signal distribution (spectral or temporal) within each frame. Naively, one can just replace the integrated signal $F(\mathbf{r},t)$ by its density $f(\mathbf{r},t,x)$ in Eq. (4) and get spectrally or temporarily resolved cumulant image:
$C_{t}^{(n)}[f(\mathbf{r},t,x)]=\sum_{k}U^{n}(\mathbf{r}-\mathbf{r}_{k})\varepsilon_{k}^{n}\varepsilon^{n}(x,\theta_{k})C_{t}^{(n)}[s_{k}(t)].$ (21)
Practically, such approach would mean that one takes the binned signal density $f(\mathbf{r},t,x_{i})$, specified for a discrete set of values $x_{i}$, and performs analysis of frame-wise temporal statistics for each bin $x_{i}$ by considering the time series $\{f(\mathbf{r},t,x_{i}\}_{t}$. However, such time series is unlikely to satisfy the requirement of having enough photons in each frame, which is important for reliable application of SOFI. (While there have already exist methods capable of resolving shot-noise induced systematic artifacts in SOFI [68], the statistical inaccuracy inevitably grows for higher-order cumulants).
Such issue can be resolved by the above-discussed “compression” of a full profile over $x$ into several scalar quantities. Following Eq. (15), one can define the signals, integrated over two spectral or temporal windows (indexed as $j=0,1$) within each frame, as follows:
[…]
where
[…]
is the $j$-th channel signal of the $k$-th emitter subjected to the environmental parameter value $\theta=\theta_{k}\equiv\theta(\mathbf{r}_{k})$.
In contrast to the binned density $f(\textbf{r},t,x_{i})$, the integrated signals $F_{j}(\mathbf{r},t)$ are intense enough to be exploited for SOFI. The cumulants can be constructed for each window separately:
$C_{j}^{(n)}(\mathbf{r})=C_{t}^{(n)}[F_{j}(\mathbf{r},t)]\\ =\sum_{k}U^{n}(\mathbf{r}-\mathbf{r}_{k})\varepsilon_{k}^{n}Y_{j}^{n}(\theta_{k})C_{t}^{(n)}[s_{k}(t)].$ (24)
In the way, inherent to SOFI, the blurring is reduced by raising the PSF $U$ in the power $n$. The expressions for the quantities $C_{0}^{(n)}(\mathbf{r})$ and $C_{1}^{(n)}(\mathbf{r})$ differ only by the factors $Y_{0}^{n}(\theta_{k})$ and $Y_{1}^{n}(\theta_{k})$ in the sums, while the leading contribution to the sum is provided by the emitters with $\mathbf{r}_{k}\approx\mathbf{r}$. Therefore, their ratio can be considered as an approximation of the local value of the compressed signal $Z(\mathbf{r})$:
$Z(\mathbf{r})\equiv\zeta(\theta(\mathbf{r}))\approx Z_{n}(\mathbf{r})\equiv\left[\frac{C_{1}^{(n)}(\mathbf{r})}{C_{0}^{(n)}(\mathbf{r})}\right]^{1/n}.$ (25)
For even $n$, the information about the sign of $Z(\mathbf{r})$ can be lost, since $Y_{j}^{n}(\theta_{k})\geq 0$ for even $n$ regardless of $\operatorname{sign}Y_{j}(\theta_{k})$. However, that limitation is typically irrelevant in practice, because non-negative weights $Q_{j}(x)\geq 0$ for integration over spectral or temporal windows ensure $Y_{j}\geq 0$.
The resulting estimator for the $n$-th order SOFI combined with sensing is
$\bar{\theta}^{(n)}(\mathbf{r})=\zeta^{-1}\left(Z_{n}(\mathbf{r})\right).$ (26)
The estimator possesses the following properties justifying its application (see Appendix B):
  1. 1.
    For a uniform (constant) parameter distribution $\theta(\mathbf{r})=\theta_{0}$, the estimator reproduces the correct parameter value for all orders $n$:
    $\bar{\theta}^{(n)}(\mathbf{r})=\theta_{0}\text{ for all {\hbox{\mathbf{r}}} and {\hbox{n}}}.$ (27)
  2. 2.
    If all the emitters are subjected to the environment parameter $\theta$ limited to certain range: $\theta_{k}\in[\theta_{\text{min}},\theta_{\text{max}}]$ for all $k$, the estimator will also belong to that range:
    $\bar{\theta}^{(n)}(\mathbf{r})\in[\theta_{\text{min}},\theta_{\text{max}}]\text{ for all {\hbox{\mathbf{r}}} and {\hbox{n}}}.$ (28)
  3. 3.
    For small variations of the parameter $\theta$ (i.e., in the linear approximation), the estimator will reproduce a weighted combination of the parameter values $\theta_{k}$ for the emitters in the vicinity of the point of interest:
    $\bar{\theta}^{(n)}(\mathbf{r})\approx\frac{\sum_{k}U^{n}(\mathbf{r}-\mathbf{r}_{k})C_{t}^{(n)}[s_{k}(t)]\varepsilon_{k}^{n}\theta_{k}}{\sum_{k}U^{n}(\mathbf{r}-\mathbf{r}_{k})C_{t}^{(n)}[s_{k}(t)]\varepsilon_{k}^{n}}.$ (29)
    For larger $n$, the estimate becomes more localized due to raising the PSF $U$ to the power $n$.
II.4 Reconstruction of parameter distribution for noisy signal
While for ideal noiseless data the constructed estimator (26) would provide reasonable local values for the investigated parameter distribution $\theta(\mathbf{r})$, its application to noisy data (especially for sparsely located emitters) deserves additional attention. The structure of expression (24) implies that the cumulant values $C_{j}^{(n)}(\mathbf{r})$ estimated from finite-sample data are more reliable near emitters ($\mathbf{r}\approx\mathbf{r}_{k}$ for certain $k$), but may contain significant relative shot noise contribution in the region between the emitters where the signals are weak. The effect becomes more pronounced for higher-order cumulants, since the effective PSF $U^{n}(\mathbf{r}-\mathbf{r}_{k})$ is narrower for larger $n$, while the contribution of shot noise is generally growing for larger cumulant orders.
The inaccuracy $\Delta\bar{\theta}^{(n)}(\mathbf{r})$ of the reconstructed parameter value $\bar{\theta}^{(n)}(\mathbf{r})$ is determined by the relative contribution of the shot noise to the quantities $C_{0}^{(n)}(\mathbf{r})$ and $C_{1}^{(n)}(\mathbf{r})$, which is, roughly, inversely proportional to the square root of the detected signal:
$\Delta\bar{\theta}^{(n)}(\mathbf{r})\propto\frac{1}{\sqrt{\left|C_{0}^{(n)}(\mathbf{r})+C_{1}^{(n)}(\mathbf{r})\right|}}.$ (30)
The likelihood of getting the estimated distribution $\bar{\theta}^{(n)}(\mathbf{r})$ for the true distribution $\theta(\mathbf{r})$ can be represented as
$L\bigl{[}\bar{\theta}^{(n)}(\mathbf{r})\big{|}\theta(\mathbf{r})\bigr{]}\propto\exp\left(-\int d^{2}\mathbf{r}\frac{\left[\bar{\theta}^{(n)}(\mathbf{r})-\theta(\mathbf{r})\right]^{2}}{2\left[\Delta\bar{\theta}^{(n)}(\mathbf{r})\right]^{2}}\right).$ (31)
In the regions with low signal, the inaccuracy $\Delta\bar{\theta}(\mathbf{r})$ can be large, making the estimate $\bar{\theta}(\mathbf{r})$ unreliable. The estimate can be refined in the spirit of Tikhonov’s regularization [69]. Very sharp features of the reconstructed distribution of the environment parameter $\theta$ from a blurred image with finite density of fluorescent emitters are likely to be artifacts rather than to provide useful information about the sample. The statement can be captured as a reduced prior probability of the distributions with large values of second derivatives:
$p_{0}[\theta(\mathbf{r})]\propto\exp\left(\frac{1}{2D^{2}}\int d^{2}\mathbf{r}\left|\nabla^{2}\theta(\mathbf{r})\right|^{2}\right),$ (32)
where the parameter $D$ represents the scale for the derivatives and $\nabla^{2}=\partial^{2}/\partial x^{2}+\partial^{2}/\partial y^{2}$ is 2-dimensional Laplacian.
According to Bayes’ formula [70, 71], the posterior probability of the true distribution $\theta(\mathbf{r})$ conditioned by the estimated distribution $\bar{\theta}^{(n)}(\mathbf{r})$ equals
$p\bigl{[}\theta(\mathbf{r})\big{|}\bar{\theta}^{(n)}(\mathbf{r})\bigr{]}\propto L\bigl{[}\bar{\theta}^{(n)}(\mathbf{r})\big{|}\theta(\mathbf{r})\bigr{]}p_{0}[\theta(\mathbf{r})].$ (33)
The distribution $\theta(\mathbf{r})$ can be reconstructed according to maximum a posteriori method:
$\theta(\mathbf{r})=\operatorname*{argmax}\limits_{\theta(\mathbf{r})}p\bigl{[}\theta(\mathbf{r})\big{|}\bar{\theta}^{(n)}(\mathbf{r})\bigr{]}.$ (34)
Practically, the parameter values are estimated at a discrete grid of points $\mathbf{r}_{ij}$. In that case, the derivatives in Eq. (33) can be replaced by finite differences, and the values $\theta(\mathbf{r}_{ij})$ can be found by unconstrained quadratic minimization (Appendix C).
II.5 Sensing method summary
To realize the proposed sensing approach with spatial super-resolution capabilities, the following procedures should be performed.
1. Model construction. The models for parameter-dependent spectral or temporal profiles $\varepsilon(x,\theta)$ of the emitters’ fluorescence and for the emitters’ blinking should be constructed (either theoretically or from empirical calibration procedure).
2. Sensing calibration. The optimal weights $Q_{0}(x)$ and $Q_{1}(x)$ of the two detection channels are defined in Appendix A. If the experimental capabilities do not allow to perform such optimal measurement, the actual weights $Q_{0}(x)$ and $Q_{1}(x)$ are defined by the available detection setup (the profiles of the channels’ sensitivity). Based on the functions $\varepsilon(x,\theta)$, $Q_{0}(x)$, and $Q_{1}(x)$, the dependence $\zeta(\theta)$ of the compressed signal on the sensed parameter should be constructed according to Eq. (16). Then, the estimator for the $n$-th order of SOFI is described by Eq. (26).
3. Data acquisition. The fluorescent response of the sample is collected in terms of 2-channel framed images $F_{k}(\mathbf{r}_{ij},t)$, where $k=0$, 1 is the channel index, $t$ is the frame index, and $\{\mathbf{r}_{ij}\}$ is the set of detection positions (pixel positions for an array detector or the scanning positions).
4. Data processing. For each channel $k$ and each detection position $\mathbf{r}_{ij}$ separately, the cumulant value $C_{k}^{(n)}(\mathbf{r}_{ij})=C_{t}^{(n)}[F_{k}(\mathbf{r}_{ij},t)]$ is calculated. For $n=2$, 3, 4 the expressions are provided by Eqs. (6)–(8) with the expectation values replaced by finite-sample averages. Alternatively, one can use cross-cumulants for improvement of signal-to-noise ratio as described in Appendix D. The local estimates of the parameter $\bar{\theta}^{(n)}(\mathbf{r})$ are calculated according to Eq. (26).
5. Data interpolation. To account for the difference of the estimate reliability at different points, the corrected parameter distribution is constructed according to Eq. (34) with the method described in Appendix C.
III Numerical modeling
III.1 Modeled setup
To illustrate the proposed combined sensing approach, we start from its application to simulated data. The modeled object includes equidistant fluorescent probes in $8\times 8$ grid with the step $d=0.266d_{\text{R}}$ (Fig. 2) or $0.532d_{R}$ (Fig. 3), where $d_{\text{R}}$ is the Rayleigh resolution limit of the microscope for classical incoherent-light imaging. Each emitter is assigned the value $\theta_{k}=\pm 1$ of the sensed parameter $\theta$ (Figs. 2(a) and 3(a)). The modeled distribution of the parameter $\theta$ has been chosen so that sensing of features with different spatial scale can be analyzed.
The blinking of the modeled emitters is described by the telegraph process with the waiting time for switching between the bright (“on”) and dark (“off”) states distributed according to the probability densities
$p_{\text{on,off}}(\Delta t)=t_{\text{on,off}}\exp\left(-\Delta t/t_{\text{on,off}}\right),$ (35)
where $t_{\text{on}}$ and $t_{\text{off}}$ are the mean durations of continuous staying the bright and dark state respectively. The natural time unit for the modeling is the SOFI frame duration. In that units, the “on” and “off” times used for the modeling are $t_{\text{on}}=2$ frames and $t_{\text{off}}=3$ frames. The total observation time is assumed to include $1.5\times 10^{4}$ frames.
We model measurement of the fluorescence signal by a $40\times 40$ pixels array detector. The pixel size is related to the distance between the emitters as $d_{\text{pix}}:d=1:5$. The number of photons, detected from each emitter per frame (in total over the whole detection plane), obeys Poisson statistics with the expectation value $10^{3}$ photons per emitter per frame. The signal is split into two spectral channels (Fig. 1(e)) in the ratio $Y_{1}:Y_{2}=0.8:1.2$ for $\theta=+1$ and $Y_{1}:Y_{2}=0.7:1.3$ for $\theta=-1$. The signal $Y_{0}$, used for normalization in denominator of Eq. (15), is defined as the sum of both channels: $Y_{0}=Y_{1}+Y_{2}$. The SOFI images are constructed according to the approach, described in Appendix D with $\sigma=3$.
We assume that the dependence of the emitters’ response to the variations of the sensed parameter $\theta$ is linear and follows Eq. (18). According to the ratios of the channel signals, defined above, the coefficients in Eq. (19) are related as $A_{01}=0$, $A_{10}=0.375A_{00}$, and $A_{11}=0.025A_{00}$. Therefore, the estimator (19) takes the form
$\bar{\theta}=\zeta^{-1}(Z)=40Z-15.$ (36)
The source code for the discussed numerical modeling is available in Supplemental Material at [URL will be inserted by publisher].
III.2 Results
Figs. 2 and 3 show the modeling results for the cases of sub-Rayleigh and relatively well-resolved imaging with sensing respectively. One can see that, as expected, the spatial resolution of the parameter inference increases for higher orders of SOFI cumulant images. To illustrate the effect, profiles corresponding to a vertical cross-section of the maps are shown in Fig. 4. For sub-Rayleigh imaging regime (Fig. 4(a)), the fine features of the parameter distribution are not visible in the results for the first and the second order cumulants, while separate minima and maxima are already present in the profile corresponding to the fourth order cumulant (i.e., the features are resolved according to Rayleigh’s criterion). For the case shown in Fig. 4(b), the features of the parameter distribution are successfully resolved for all the considered cumulant orders. However, the accuracy of the parameter inference is much higher for the fourth order of SOFI: the mean squared error (MSE), defined as $\operatorname{MSE}(\bar{\theta})=\langle(\bar{\theta}(\mathbf{r}_{k})-\theta_{k})^{2}\rangle$ takes the values 0.40, 0.11, and 0.003 for the first, the second, and the fourth order of SOFI cumulants respectively.
Figure 2: Modeling results for sub-Rayleigh imaging with sensing. (a) The object is modeled as a grid of emitters with the step $d=0.266d_{\text{R}}$. The emitters are assigned the parameter values $\theta=+1$ (red dots) and $\theta=-1$ (blue dots). The vertical dashed line shows the cross-section used for extracting the profiles in Fig. 4. (b-d) The reconstructed distribution $\bar{\theta}^{(n)}(\mathbf{r})$ of the sensed parameter for $n=1$ (b), 2 (c), 4 (d). The color map is the same for all plots (dark blue and light yellow correspond to $\bar{\theta}=-1$ and $\bar{\theta}=+1$ respectively). The spatial coordinates $x$ and $y$ are normalized by the Rayleigh limit $d_{\text{R}}$.
Figure 3: Modeling results for relatively well-resolved imaging with sensing. The step size of emitters’ grid is $d=0.532d_{\text{R}}$. Other parameters and notations are the same as in Fig. 2.
Figure 4: Inferred parameter distributions for the vertical cross-section of the maps, shown in Figs. 2 and 3. Dot-dashed, dashed, and solid lines correspond to $n=1$, 2, and 4 respectively. Blue and red dots indicate the true values of the parameter $\theta_{k}$ ($-1$ and $+1$ respectively) for the emitters along the considered sample cross-section.
IV Model experiment
IV.1 Setup description
To get closer to real-world verification of the proposed technique, we also performed the following model experiment. We used a paper screen as a model of fluorescent sample. The screen was irradiated with 13 KY-008 lasers controlled by the microcontroller board Arduino MEGA [72] and capable of independent switching between “dark” and “bright” states. The image of the screen was recorded with Thorlabs DCC1545M-GL camera with 50 mm lens (Fig. 5). Switching of lasers was synchronized to the frame rate of the camera (24 fps, approximately 42 ms per frame). To emulate the diffraction limit, the lens was placed 2 mm closer to the camera than it was necessary to get a clear picture. The resulting image of unfocused laser spots is shown in Fig. 6. Each laser spot was used for modeling a fluorescent particle (temperature nanosensor).
Figure 5: The scheme of the model experiment. The lasers controlled with Arduino microcontroller board irradiate the paper screen. The light scattered by the screen is collected with lens to the Thorlabs camera.
To model fluorescent particles with different temperature, the lasers were divided into “hot” ($\theta=+1$) and “cold” ($\theta=-1$) ones. The parameter value $\theta$ (normalized “temperature”) was encoded in the “relaxation time” — the time between “excitation pulse” and “fluorescent emission”. To denote “excitation pulse” during data recording, all lasers blinked twice (i.e., a series of 2 one-frame flashes of lasers separated by 1 frame of darkness was produced). The “fluorescent emission” was 1-frame flashing of a laser. For the “hot” lasers, the “relaxation time” was 5 frames with standard deviation of 2 frames, while the “cold” lasers had “relaxation time” of 10 frames with the same standard deviation.
Figure 6: Image of blurred laser spots on the paper screen for the lens being placed 2 mm closer to the camera than it is necessary for a clear image. Numbers are used for indexing the laser beams during further analysis.
IV.2 Data acquisition
The data acquisition was based on 3 time scales (Fig. 7):
  • Camera frame (42 ms) represented the natural elementary unit of time.
  • “Excitation cycle” lasted 20 camera frames and included an “excitation pulse” and “fluorescent emission” of each active laser.
  • SOFI frame included 20 “excitation cycles” (400 camera frames) and determined blinking rate of the modeled fluorescent particles.
For each SOFI frame, each laser was randomly assigned a “bright” (active) or “dark” (inactive) state to model stochastic blinking. The probabilities of the two states were equal to 1/2, and the sampling was performed independently for all lasers and all SOFI frames. For each “excitation cycle” withing a SOFI frame, each laser being in a “bright” state was randomly assigned its relaxation time (delay of “fluorescent emission”). The delay was sampled according to normal distribution with the mean determined by the model temperature of the laser (“hot” or “cold”). The random sampling was performed by the Arduino microcontroller. Then, the microcontroller produced an “excitation pulse” and made each active laser to flash after the time delay assigned to it. After modeling “fluorescent emission” of all the active lasers withing the current cycle (when all the lasers are switched off), the microcontroller waited for 1 camera frame, and repeated the cycle. In the end, we got a video containing intensity distribution in time and space. The total acquisition process included $2\times 10^{5}$ camera frames ($10^{4}$ excitation cycles, and 500 SOFI frames).
IV.3 Data processing and results
First, the recorded video was divided into frames with ffmpeg program [73] and processed with Python script. The script calculated total intensity for each camera frame and identified “excitation pulses” according to specific alternation of high and low intensity of frames. The frames designating the “excitation pulse” were excluded from further calculations. The resulting series of “excitation cycles” was split into groups of 20 cycles, each group forming a SOFI frame. According to the notations, introduced in Section II, the SOFI frames are indexed by $t=1$, …, $N_{\text{f}}$.
The signal for each “excitation cycle” was split into two channels and accumulated pixelwise. Each SOFI frame accumulates the signals from the “excitation cycles” included in that frame. The “hot” signal $Y_{1}(i,j;t)$ is the pixelwise sum of intensities for the 1st - 7th camera frames after the $t$-th “excitation pulse” accumulated for 20 cycles, where indices $i$ and $j$ enumerate pixel positions. The “cold” signal $Y_{2}(i,j;t)$ is defined in the same way starting from the 8th camera frame after the “excitation pulse” (Fig. 7).
Figure 7: Data acquisition timeline. (a) A SOFI frame consist of 20 “excitation cycles”. (b) An example of “excitation cycle” timeline, where two pulses appear in the “hot” time window and two pulses - in the “cold” time window.
The intensity (mean signal value) and the 2nd, 3rd, and 4th order cumulants for the two channels where defined according to Eqs. (6)–(9). Here, only zero time lag autocumulants were used since the signal did not suffer from significant shot noise. The obtained intensity and cumulant images are shown in Fig. 8. For visualization purposes, the “cold” and “hot” channels were assigned to the blue and red channels of RGB palette respectively. Since the cumulants of the 3rd and the 4th order can be negative, the absolute values of the cumulants were used to create the image. One can see that the intensity distribution (Fig. 8(a)) is blurred like original unfocused image (Fig. 6(b)). However, it is possible to distinguish the light of “hot” and “cold” lasers because of different relaxation times. The light spots are more visible in the 2nd order cumulant image (Fig. 8(b)), but the effect of blurring is still strong. The 3rd and the 4th order cumulant images (Fig. 8(c,d)) have the best quality and allow one to distinguish clearly spots of “hot” and “cold” lasers. To summarize, the experiment is in a good agreement with the theory, and the calculation of cumulants of intensity fluctuations allows one to visualize the particles with different relaxation times.
Figure 8: The two-channel signal for intensity (a) and cumulants of the 2nd (b), 3rd (c), and 4th (d) order.
IV.4 Reconstruction of local “temperature”
To fit the above-described model experiment into the theoretical framework, discussed in Section II, we assign the normalized dimensionless “temperature” equal to $\theta=-1$ to “cold” lasers and $\theta=+1$ to “hot” lasers. The emission profile variable $x$ corresponds to the camera frame number within the “excitation cycle” in the considered case. For such definitions, the model emission profile is described as
$\varepsilon(x,\theta)=\frac{1}{\sqrt{2\pi}\sigma}\exp\left(-\frac{(x-7.5+2.5\theta)^{2}}{2\sigma^{2}}\right),$ (37)
where $\sigma=2$.
A simple threshold filter is used to split the two detection channels (see Fig. 1(f)):
$Y_{1}=\sum_{x\leq 7}\varepsilon(x,\theta),\quad Y_{2}=\sum_{x\geq 8}\varepsilon(x,\theta),\quad Y_{0}=Y_{1}+Y_{2}.$ (38)
The function $\zeta(\theta)$, calculated according to Eq. (16), is plotted in Fig. 9. The estimators $\bar{\theta}^{(n)}(\mathbf{r})$ are defined by Eq. (26). Further, the reconstructed distribution can be refined using the method outlined in the Section II.4 and Appendix C.
Figure 9: Relation between the compressed signal $Z=\zeta(\theta)$ and the sensed parameter $\theta$ (normalized dimensionless “temperature”) for the model experiment, calculated according to Eqs. (16), (37), and (38).
The results are presented in Figs. 10 and 11. As predicted, usage of higher-order cumulants enhances spatial resolution and enables analysis of finer features of the sample. According to the standard performance of SOFI, using fourth order cumulant instead of intensity leads to approximately twice narrower efficient PSF.
Figure 10: Reconstructed distribution of the model normalized temperature $\theta$ for analysis of intensity (a), the 2nd (b), 3rd (c), and 4th (d) order cumulants. Black lines indicate the border of the region where non-zero signal was collected; outside that region the temperature values are effectively extrapolated by the approach presented in Subsection II.4. Black dots shows the central positions of the laser spots (positions of the modeled fluorescent emitters).
Figure 11: Cross-section of the reconstructed distributions of the model normalized temperature $\theta$, shown in Fig. 10, for the vertical line passing through the emitter no. 3. Black markers shows true values of the parameter $\theta$ for the emitters no. 1, 3, 5, and 9 adjacent to the considered cross-section.
IV.5 Discussion
To get quantitative estimates of the proposed approach performance, we compared the estimated parameter values $\bar{\theta}_{k}=\bar{\theta}(\mathbf{r}_{k})$ at the positions of the modeled fluorescent particles (centers of the laser spots) with the true values $\theta_{k}=\pm 1$ (Fig. 12(a)). Fig. 12(b) shows the dependence of the mean squared error (MSE), defined as $\operatorname{MSE}(\theta)=\langle(\bar{\theta}_{k}-\theta_{k})^{2}\rangle_{k}$, on the cumulant order $n$. The results indicate that the least accurate estimate corresponds to the blurred distribution obtained for the traditional intensity-based approach. The estimation accuracy is saturated for $n=3$, while for $n=4$ the error even increases slightly because of stronger artifacts and noise enhancement typical for higher-order cumulants. Such saturation is explained by the specific profile of blurring (containing a plateau, as one can see from Fig. 11) — see Appendix E for details.
Figure 12: Performance of the proposed approach. (a) Comparison of estimated parameter values $\bar{\theta}_{k}=\bar{\theta}(\mathbf{r})_{k})$ at the positions of the modeled fluorescent particles (centers of the laser spots) with the true values $\theta_{k}=\pm 1$ for different orders of cumulants used for the analysis (blue circles, yellow squares, green diamond makers, and red stars correspond to $n=1$, 2, 3, and 4 respectively). (b) Mean squared error of the estimation as the function of the cumulant order $n$.
V Conclusions
We proposed a simple and efficient technique capable of combining sensing with super-resolution imaging, based on non-trivial temporal statistics typical for blinking fluorescent nanoemitters. The approach is universal since the dependence of the fluorophores response on the characteristic to be reconstructed can take different form: changes of the spectrum, fluorescence lifetime, etc. It is shown that splitting the fluorescence signal into just two detection channels at each spatial detection point is sufficient for reliable reconstruction of the local characteristic distribution.
The conducted numerical simulations and model experiment proof soundness and efficiency of the approach and can be used as a guidance for the real-world experiments with, for example, living cells and structured fluorescent nanodiamonds [74] or blinking graphene quantum dots ensembles [75]. In agreement with the theoretical predictions, usage of cumulants instead of emission intensity enhances sharpness of the reconstructed parameter distribution, enables analysis of smaller features of the sample, and ensures higher accuracy of the estimation.
Proposed methodology will be useful for local measurements of living cells’ physical parameters, such as temperature, pH, electric field gradient, presence of reactive oxygen spices, i.e. the characteristics which could be recognized as spectral changes of appropriate fluorophore delivered targetly to the wanted cell or even organelle. In its turn, this advanced local sensing tool could give an important insights into fundamental life science, which still lacks instruments capable of monitoring the cellular balance with nanometer spatial resolution to guarantee measurements consistency for that dynamic system.
We believe that the proposed approach and its proof-of-concept experiment can pave the way for further research, including the following possible directions:
  • detailed analysis and optimization of the combined sensing and imaging approach in terms of classical and quantum Fisher information;
  • reconstruction of the sample characteristic distribution by combined usage of several orders of cumulants simultaneously;
  • extension of the approach to multimodal sensing, when several characteristics of the sample are reconstructed simultaneously;
  • implementation of Fourier spectrum re-weighting for additional de-blurring and cumulant orders combination;
  • application of the “sliding window” approach [50] for more efficient processing of many-pixel images.
Acknowledgements.
This study was accomplished through the financial support of Research Council of Finland (Flagship Programme PREIN, decision 346518; research project, decision 368653); QuantERA project EXTRASENS, in Finland funded through Research Council of Finland, decision 361115), Horizon Europe MSCA FLORIN Project 101086142.
Appendix A Locally optimal signal compression
Here, we show that the information about the environment parameter $\theta$, contained in the normalized signal density profile $g(x)\equiv\tilde{f}(\mathbf{r}_{0},t,x)$, defined by Eq. (14), can be efficiently compressed into a ratio $Z$ of two scalar characteristics $Y_{i}$ (Eq. (15)) containing the same information. We assume that the signal profile is measured at a discrete set of bins $\{x\}$ and contains shot noise $\delta(x)$ satisfying Poisson statistics (i.e., we analyze shot noise for a fixed blinking realization since the blinking influences the in-frame signal profile as a whole):
$g(x)\mapsto g_{\text{noisy}}(x)=g(x)+\delta(x),$ (39)
where $\operatorname{E}[\delta(x)]=0$, $\operatorname{E}[\delta(x)\delta(x^{\prime})]\approx g(x)\delta_{xx^{\prime}}$. The high signal-to-noise ratio (SNR) regime is considered, i.e., $|\delta(x)|\ll|g(x)|$.
The Fisher information [76] of the normalized density profile $g(x)$ relatively to the parameter $\theta$ can be calculated as
$F_{\text{full}}(\theta)=\sum_{x}\frac{\left[g^{\prime}(x)\right]^{2}}{g(x)}-\frac{1}{\sum_{x}g(x)}\Bigl{(}\sum_{x}g^{\prime}(x)\Bigr{)}^{2},$ (40)
where the second term describes “information loss” caused by the normalization [77] and $g^{\prime}(x)=dg(x)/d\theta$.
Substitution of $g_{\text{noisy}}(x)$ into Eq. (15) and replacement of the integrals by sums yields
$Z\mapsto Z_{\text{noisy}}\\ \approx Z\left(1+\frac{\sum_{x}\delta(x)Q_{1}(x)}{\sum_{x}g(x)Q_{1}(x)}-\frac{\sum_{x}\delta(x)Q_{0}(x)}{\sum_{x}g(x)Q_{0}(x)}\right).$ (41)
In the considered high-SNR regime, $Z_{\text{noisy}}$ represents a normally distributed random variable with the expectation value $\operatorname{E}(Z_{\text{noisy}})=Z$ and the variance
$\operatorname{Var}(Z_{\text{noisy}})=\frac{1}{Y_{0}^{2}}\sum_{x}g(x)\bigl{[}Q_{1}(x)-Q_{0}(x)Z\bigr{]}^{2}.$ (42)
Therefore, Fisher information for the signal $Z_{\text{noisy}}$ relatively to the sensed parameter $\theta$ equals (see, e.g., Ref. [77])
$F_{Z}(\theta)=\frac{1}{\operatorname{Var}(Z_{\text{noisy}})}\left(\frac{dZ}{d\theta}\right)^{2}.$ (43)
According to general properties of Fisher information [78, 79], the compressed signal $Z_{\text{noisy}}$ cannot contain more information than the initial normalized profile $g_{\text{noisy}}(x)$ for any choice of the weights $Q_{0}(x)$ and $Q_{1}(x)$:
$F_{Z}(\theta)\leq F_{\text{full}}(\theta).$ (44)
On the other hand, by a specific choice of those weights, one can achieve equality of the two quantities. One can define
$Q_{0}(x)=\frac{q_{0}(x)}{\sum_{x^{\prime}}q_{0}(x^{\prime})g(x^{\prime})},\;Q_{1}(x)=\frac{q_{1}(x)}{\sum_{x^{\prime}}q_{1}(x^{\prime})g^{\prime}(x^{\prime})},$ (45)
where
$q_{0}(x)=g(x)-g^{\prime}(x)\frac{\sum_{x^{\prime}}g(x^{\prime})g^{\prime}(x^{\prime})}{\sum_{x^{\prime}}\left[g^{\prime}(x^{\prime})\right]^{2}}$ (46)
and
$q_{1}(x)=\frac{g^{\prime}(x)}{g(x)}-\frac{\sum_{x^{\prime}}g^{\prime}(x^{\prime})}{\sum_{x^{\prime}}g(x^{\prime})}.$ (47)
That specific choice of the weights ensures that
$Y_{0}=1,\quad Y_{1}=0,\quad\frac{dY_{0}}{d\theta}=0,\quad\frac{dY_{1}}{d\theta}=1$ (48)
and
$Z=0,\quad\frac{dZ}{d\theta}=1$ (49)
at the considered value of the parameter $\theta$ (the one used for construction of the weights). Substituting Eqs. (48) and (49) into Eqs. (42) and (43), we get the results $\operatorname{Var}(Z_{\text{noisy}})=1/F_{\text{full}}(\theta)$ and
$F_{Z}(\theta)=F_{\text{full}}(\theta).$ (50)
The choice of the weights is locally optimal. Generally, for values of $\theta$, other than the specific one used for the weights construction, the equality (50) may be violated. It is worth noting that in the high-SNR limit and in the vicinity of the considered parameter value $\theta$ the estimator
$\bar{\theta}=\zeta^{-1}(Z)\approx\theta+Z$ (51)
with the weights described by Eq. (45) coincidence with the locally unbiased estimator, saturating Cramer-Rao bound [80].
Appendix B General properties of cumulant-based estimator for noiseless data
To be physically reasonable, the estimator $\bar{\theta}^{(n)}(\mathbf{r})$, defined by Eq. (26) should satisfy several general requirements when applied to noiseless data. The requirements are provided by Eqs. (27)–(29). Here, we proof those statements.
1. For a spatially uniform distribution $\theta(\mathbf{r})=\theta_{0}$, one has $Y_{j}(\theta_{k})=Y_{j}(\theta_{0})$ for all emitter indices $k$. Therefore, according to Eq. (24),
$\frac{C_{1}^{(n)}(\mathbf{r})}{C_{0}^{(n)}(\mathbf{r})}=\frac{Y_{1}^{n}(\theta_{0})}{Y_{0}^{n}(\theta_{0})}=\zeta^{n}(\theta_{0})\text{ for all {\hbox{\mathbf{r}}}},$ (52)
and
$Z_{n}(\mathbf{r})=\zeta(\theta_{0})\quad\Rightarrow\quad\bar{\theta}^{(n)}(\mathbf{r})=\theta_{0}.$ (53)
2. If the parameter values $\theta_{k}$ in the vicinity of the emitters are bounded as $\theta_{k}\in[\theta_{\text{min}},\theta_{\text{max}}]$, the monotonicity assumption implies that the quantities $\zeta(\theta_{k})$ are also bounded as $\zeta(\theta_{k})\in[\zeta_{\text{min}},\zeta_{\text{max}}]$, where $\zeta_{\text{min}}=\min(\zeta(\theta_{\text{min}}),\zeta(\theta_{\text{max}}))$ and $\zeta_{\text{max}}=\max(\zeta(\theta_{\text{min}}),\zeta(\theta_{\text{max}}))$. Under the assumption that $\zeta_{\text{min}}\geq 0$ and $U(\mathbf{r}-\mathbf{r}_{k})\geq 0$, for each emitter, the following relation holds:
$Y_{1}^{n}(\theta_{k})\in Y_{0}^{n}(\theta_{k})\times[\zeta_{\text{min}}^{n},\zeta_{\text{max}}^{n}].$ (54)
Therefore, for any $\mathbf{r}$,
$C_{1}^{(n)}(\theta_{k})\in C_{0}^{(n)}(\theta_{k})\times[\zeta_{\text{min}}^{n},\zeta_{\text{max}}^{n}].$ (55)
Finally, $Z_{n}(\mathbf{r})\in[\zeta_{\text{min}},\zeta_{\text{max}}]$ and, therefore, $\bar{\theta}^{(n)}(\mathbf{r})\in[\theta_{\text{min}},\theta_{\text{max}}]$.
3. If the variations $\delta\theta_{k}=\theta_{k}-\theta_{0}$ are small, one can use the linear approximation and get
$Y_{j}^{n}(\theta_{k})\approx Y_{j}^{n}(\theta_{0})+nY_{j}^{n-1}(\theta_{0})Y_{j}^{\prime}(\theta_{0})\delta\theta_{k},$ (56)
[…]
$C_{j}^{(n)}(\mathbf{r})\approx Y_{j}^{n}(\theta_{0})\sum_{k}U^{n}(\mathbf{r}-\mathbf{r}_{k})\varepsilon_{k}^{n}C_{t}^{(n)}[s_{k}(t)]\\ +nY_{j}^{n-1}(\theta_{0})Y_{j}^{\prime}(\theta_{0})\sum_{k}U^{n}(\mathbf{r}-\mathbf{r}_{k})\varepsilon_{k}^{n}C_{t}^{(n)}[s_{k}(t)]\delta\theta_{k}$ (57)
and
$Z_{n}(\mathbf{r})\approx\zeta(\theta_{0})\Biggl{[}1+\left(\frac{Y_{1}^{\prime}(\theta_{0})}{Y_{1}(\theta_{0})}-\frac{Y_{0}^{\prime}(\theta_{0})}{Y_{0}(\theta_{0})}\right)\\ \times\frac{\sum_{k}U^{n}(\mathbf{r}-\mathbf{r}_{k})\varepsilon_{k}^{n}C_{t}^{(n)}[s_{k}(t)]\delta\theta_{k}}{\sum_{k}U^{n}(\mathbf{r}-\mathbf{r}_{k})\varepsilon_{k}^{n}C_{t}^{(n)}[s_{k}(t)]}\Biggr{]}$ (58)
Since
$\zeta(\theta_{0}+\delta\theta)\approx\zeta(\theta_{0})\left[1+\left(\frac{Y_{1}^{\prime}(\theta_{0})}{Y_{1}(\theta_{0})}-\frac{Y_{0}^{\prime}(\theta_{0})}{Y_{0}(\theta_{0})}\right)\delta\theta\right],$ (59)
one can easily find the solution of the equation $\zeta(\theta)=Z_{n}(\mathbf{r})$ in the linear approximation and conclude that
$\bar{\theta}^{(n)}(\mathbf{r})=\theta_{0}+\frac{\sum_{k}U^{n}(\mathbf{r}-\mathbf{r}_{k})\varepsilon_{k}^{n}C_{t}^{(n)}[s_{k}(t)]\delta\theta_{k}}{\sum_{k}U^{n}(\mathbf{r}-\mathbf{r}_{k})\varepsilon_{k}^{n}C_{t}^{(n)}[s_{k}(t)]}\\ =\frac{\sum_{k}U^{n}(\mathbf{r}-\mathbf{r}_{k})\varepsilon_{k}^{n}C_{t}^{(n)}[s_{k}(t)]\theta_{k}}{\sum_{k}U^{n}(\mathbf{r}-\mathbf{r}_{k})\varepsilon_{k}^{n}C_{t}^{(n)}[s_{k}(t)]}.$ (60)
Appendix C Expressions for regularized parameter estimation
For parameter values, defined at a regular grid $\mathbf{r}_{ij}$, the derivatives in Eq. (32) can be approximated as finite differences. Then, the expression for log-probability takes the form
$-\log p\left[\theta(\mathbf{r}_{ij})|\bar{\theta}^{(n)}(\mathbf{r}_{ij})\right]\propto\sum_{i,j}\frac{[\bar{\theta}^{(n)}(\mathbf{r}_{ij})-\theta(\mathbf{r}_{ij})]^{2}}{2\left[\Delta\bar{\theta}^{(n)}(\mathbf{r}_{ij})\right]^{2}}\\ +\sum_{i,j}\frac{[2\theta(\mathbf{r}_{ij})-\theta(\mathbf{r}_{i-1,j})-\theta(\mathbf{r}_{i+1,j})]^{2}}{2D^{2}}\\ +\sum_{i,j}\frac{[2\theta(\mathbf{r}_{ij})-\theta(\mathbf{r}_{i,j-1})-\theta(\mathbf{r}_{i,j+1})]^{2}}{2D^{2}}.$ (61)
The sums representing the second and the third terms do not include boundary values of $i$ and $j$ respectively. The parameter $D$ scales as $D=\text{const}\times\langle\Delta\bar{\theta}^{(n)}(\mathbf{r}_{ij})\rangle w$, where $w$ is the PSF width expressed in the grid steps and representing the expected minimal size of the reconstructed distribution features, which are not artifacts yet. The proportionality coefficient can be calculated from detailed theoretical investigation of the noise models, but, in practice, it can be just chosen empirically.
Unconstrained minimization of the quantity, defined by the right-hand side of Eq. (61), is trivial and consists in linear inversion according to the least-squares method. One of the approaches to obtaining physically reasonable constrained estimates $\theta_{\text{min}}\leq\theta(\mathbf{r}_{ij})\leq\theta_{\text{max}}$ is to perform constrained quadratic optimization, which, however, becomes much more computationally hard for a large number of grid points $\mathbf{r}_{ij}$. A more efficient way to get reasonable results is to constrain the initial estimate $\bar{\theta}^{(n)}(\textbf{r}_{ij})$ to the range $[\theta_{\text{min}},\theta_{\text{max}}]$ before applying the minimization of the expression (61). If all the values satisfy the constraint $\bar{\theta}^{(n)}(\textbf{r}_{ij})\in[\theta_{\text{min}},\theta_{\text{max}}]$, the minimizing set $\{\theta(\textbf{r}_{ij}\}$ will also belong to the physical range with certainty.
Appendix D SOFI with cross-cumulants
The expression (10) for a the zero time lag cross-cumulant
$C_{t}^{(n)}[F(\mathbf{r}^{(1)},t),\ldots,F(\mathbf{r}^{(n)},t)]\\ \propto\sum_{k}U^{n}(\bar{\mathbf{r}}-\mathbf{r}_{k})\varepsilon_{k}^{n}C_{t}^{(n)}[s_{k}(t)],$ (62)
coincides with Eq. (4) for an autocumulant $C_{t}^{(n)}[F(\bar{\mathbf{r}},t)]$ up to a constant multiplier. Therefore, to improve SNR, one can replace the autocumulant signal $C_{t}^{(n)}(\mathbf{r})$ used in SOFI by a weighted combination of cross-cumulants:
$C_{t}^{(n)}(\mathbf{r})\mapsto\tilde{C}_{t}^{(n)}(\mathbf{r})=\sum_{\begin{array}[]{c}\scriptstyle\mathbf{r}^{(1)},\ldots,\mathbf{r}^{(n)}:\\ \scriptstyle\mathbf{r}^{(1)}+\cdots+\mathbf{r}^{(n)}=n\mathbf{r}\end{array}}w_{\mathbf{r}^{(1)},\ldots,\mathbf{r}^{(n)}}\\ \times C_{t}^{(n)}[F(\mathbf{r}^{(1)},t),\ldots,F(\mathbf{r}^{(n)},t)].$ (63)
During processing of the simulated data, we choose the weights as
$w_{\mathbf{r}^{(1)},\ldots,\mathbf{r}^{(n)}}\propto\exp\left(-\frac{\sum_{j}\left|\mathbf{r}^{(j)}\right|^{2}}{\sigma^{2}}\right)$ (64)
with $\sigma=3\delta$, where $\delta$ is the detection step (detector pixel size, recalculated to the object plane of the microscope). As a compromise between noise reduction and computational complexity, we additionally restrict the sum in Eq. (63) by the condition $\sum_{j}\left|\mathbf{r}^{(j)}\right|^{2}\leq 5\sigma^{2}$.
The approach can also be extended to non-zero time lag cross-cumulants, but we do not consider such modification since the goal of simulations is illustration of the proposed sensing approach rather than achievement of the ultimate accuracy of SOFI results.
The source code of the described method implementation is available in Supplemental Material at [URL will be inserted by publisher].
Appendix E Influence of PSF shape on resolution enhancement
According to Eq. (4), the resolution enhancement in SOFI is provided by raising PSF $U(\Delta\mathbf{r})$ to the power $n$ for the $n$-th order cumulant image. If the PSF can be approximated by a Gaussian profile,
$U(\Delta\mathbf{r})\propto\exp\left(-\frac{\Delta\mathbf{r}^{2}}{2w^{2}}\right),$ (65)
it preserves the shape after raising to the power $n$ and acquires $\sqrt{n}$ times smaller width (Fig. 13(a)):
$U^{n}(\Delta\mathbf{r})\propto\exp\left(-n\frac{\Delta\mathbf{r}^{2}}{2w^{2}}\right)=\exp\left(-\frac{\Delta\mathbf{r}^{2}}{2w_{n}^{2}}\right),$ (66)
where $w_{n}=w/\sqrt{n}$. Similar scaling is characteristic for the Airy disk profile (so called “sombrero” function [81]), obtained for the blurring caused by diffraction on a lens with finite-radius aperture (Fig. 13(b)). Such behavior of typical PSF shapes allows one to state that $n$-th order SOFI enhances resolution by the factor $\sqrt{n}$.
However, the behavior can be significantly different for other PSF shapes. In the limiting case of PSF being the rectangular function
$U(\Delta\mathbf{r})\propto\operatorname{rect}(|\Delta\mathbf{r}|/w)=\left[\begin{array}[]{cc}1&|\Delta\mathbf{r}|<w,\\ 0&\text{otherwise},\end{array}\right.$ (67)
raising it to any power $n>0$ will not influence the blurring width (Fig. 13(c)):
$U^{n}(\Delta\mathbf{r})\propto U(\Delta\mathbf{r}).$ (68)
In the intermediate case of convolved Gaussian and rectangular profiles,
$U(\Delta\mathbf{r})\propto\exp\left(-\frac{\Delta\mathbf{r}^{2}}{2w_{\text{G}}^{2}}\right)*\operatorname{rect}(|\Delta\mathbf{r}|/w_{\text{R}}),$ (69)
the resolution enhancement scales with the growth of $n$ much slower than $\sqrt{n}$ (almost saturates, Fig. 13(d, e)).
Figure 13: Influence of PSF shape on resolution enhancement. (a-d) Profiles of PSF (with argument normalized by full width at half maximum — FWHM) raised to the power $n=1$ (blue solid lines), 2 (yellow dashed lines), 4 (greed dotted lines), and 8 (red dot-dashed lines) for Gaussian (a), “sombrero” (b), rectangular (c) profiles, and the convolution of Gaussian and rectangular profiles (d) with $w_{\text{G}}:w_{\text{R}}=1:10$. (e) Normalized FWHM of PSF raised to the power $n$ for for Gaussian (blue solid lines), “sombrero” (yellow dashed lines), rectangular (greed dotted lines) profiles, and the convolution of Gaussian and rectangular profiles (red dot-dashed lines).